# “How Big is Big Enough?” Adjusting Model Size in Continual Gaussian Processes

### v1

@@ 3.3 Online Sparse Gaussian Processes
^$\widehat{\mathcal{L}}=\mathbb{E}_{q_{n}(f)}\left[\log p(\mathbf{y}_{n}|f)\right]+\mathbb{E}_{q_{n}(f)}\left[\log l(\mathbf{a})\right]-\operatorname{KL}\left[q_{n}(\mathbf{b})\mid\mid p(\mathbf{b}|\theta_{n})\right]-\log\mathcal{N}(\mathbf{a};0,\mathbf{K}^{\prime}_{\mathbf{aa}}+\mathbf{D}_{\mathbf{a}})\,,$ (7)
@@
Although $\widehat{\mathcal{L}}$ is a useful training objective for general likelihoods, the regression case we consider allows us to analytically find $q(\mathbf{b})$ [2] (see Appendix G) resulting in the lower bound
^$\widehat{\mathcal{L}}=\log\mathcal{N}\left(\hat{\mathbf{y}};\mathbf{0},\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b}\hat{\mathbf{f}}}\!+\!\Sigma_{\hat{\mathbf{y}}}\right)\!+\!\Delta_{\mathbf{a}}\!-\!\frac{1}{2}\operatorname{tr}\left[\mathbf{D}_{\mathbf{a}}^{-1}(\mathbf{K}_{\mathbf{aa}}\!-\!\mathbf{Q}_{\mathbf{aa}})\right]\!-\!\frac{1}{2\sigma^{2}}\operatorname{tr}(\mathbf{K}_{\mathbf{ff}}\!-\!\mathbf{Q}_{\mathbf{ff}})\,,$ (8)
$\hat{\mathbf{y}}=\left[\begin{array}[]{c}\mathbf{y}_{n}\\ \mathbf{D}_{\mathbf{a}}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}}\end{array}\right],\mathbf{K}_{\hat{\mathbf{f}}\mathrm{b}}=\left[\begin{array}[]{l}\mathbf{K}_{\mathbf{fb}}\\ \mathbf{K}_{\mathbf{ab}}\end{array}\right],\Sigma_{\hat{\mathbf{y}}}=\left[\begin{array}[]{cc}\sigma_{\hat{y}}^{2}\mathrm{I}&\mathbf{0}\\ \mathbf{0}&\mathbf{D}_{\mathbf{a}}\end{array}\right],$ (15)
$\quad\Delta_{\mathbf{a}}=-\frac{1}{2}\log\frac{|\mathbf{S}_{\mathbf{a}}|}{|\mathbf{K}_{\mathbf{aa}}^{\prime}||\mathbf{D}_{\mathbf{a}}|}+\frac{M_{\mathbf{a}}}{2}\log(2\pi)-\frac{1}{2}\mathbf{m}_{\mathbf{a}}^{T}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}}+\frac{1}{2}\mathbf{m}_{\mathbf{a}}^{T}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{D}_{\mathbf{a}}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}}\,,$ (16)
with $\mathbf{Q}_{\mathbf{ff}}=\mathbf{K}_{\mathbf{fb}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{bf}}$ and $\mathbf{Q}_{\mathbf{aa}}=\mathbf{K}_{\mathbf{ab}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{ba}}$. All covariances are computed using the new hyperparameters $\theta_{n}$, except for $\mathbf{K}_{\mathbf{aa}}^{\prime}$ which is the covariance for the prior distribution $p(\mathbf{a}\mid\theta_{o})$. Finally, $M_{\mathrm{a}}=|\mathbf{a}|$ is the number of inducing points used at the previous batch. The computational complexity and memory requirements for calculating $\widehat{\mathcal{L}}$ at each batch is $O(N_{n}M_{\mathbf{b}}^{2}+M_{\mathbf{b}}^{3})$ and $O(M_{\mathbf{b}}^{2})$ respectively where $M_{\mathbf{b}}$ is the total number of inducing points for the current batch.
To create a fully black-box solution, we still need to specify how to select the hyperparameters $\theta_{n}$, the number of inducing variables $M_{\mathbf{b}}$, and the inducing inputs $\mathbf{Z}_{n}$. We will always select $\theta_{n}$ by maximising $\widehat{\mathcal{L}}$ using L-BFGS. To select the locations $\mathbf{Z}_{n}$, we use the “greedy variance” criterion [12, 13, 3] . This leaves only the number of inducing variables $M_{\mathbf{b}}$ to be selected.
@@ 4 Automatically Adapting Approximation Capacity
We propose a method for adjusting the capacity of the approximation $M_{\bf b}$ to maintain accuracy. We propose to keep inducing points from old batches fixed, and select new inducing points from each incoming batch, with their locations set using the “greedy variance” criterion [3, 12, 13]. While optimising all inducing points does lead to a strictly better approximation, we avoid this for the sake of simplicity. The question remains: To achieve a level of accuracy, “how big is big enough?” To answer this, we consider the online ELBO as a function of the capacity ^$\widehat{\mathcal{L}}(M_{\bf b})$, and propose a threshold after which to stop adding new inducing variables.
@@ 4.1 Online Log Marginal Likelihood (LML) Upper Bound
The problem of selecting a sufficient number of inducing variables is also still open in the batch setting. One possible strategy is to use an upper bound to the marginal likelihood [42] to bound $\mathrm{KL}[q(f)||p(f|\mathbf{y})]\leq\mathcal{U}-\mathcal{L}$, and stop adding inducing variables once this is below a tolerance $\alpha$. To extend this strategy to online learning, we begin by deriving an online upper bound, as a counterpart to the online ELBO from eq. (8). We follow the same strategy as Titsias [42], by considering the highest possible value that our lower bound can attain. While in full-batch inference this is equal to the true LML, in our case this is obtained by keeping the inducing inputs from the previous iteration, and adding each new datapoint to the inducing set:
:=^$\mathcal{L}^{*}:=\widehat{\mathcal{L}}(N_{n}+M_{\bf a})=\log\mathcal{N}\left(\hat{\mathbf{y}};\,\mathbf{0}\,,\mathbf{K}_{\hat{\mathbf{f}}\hat{\mathbf{f}}}+\Sigma_{\hat{y}}\right)+\Delta_{\mathbf{a}}\quad\text{with}\quad\mathbf{K}_{\hat{\mathbf{f}}\hat{\mathbf{f}}}=\left[\begin{array}[]{cc}\mathbf{K}_{\mathbf{ff}}&\mathbf{K}_{\mathbf{fa}}\\ \mathbf{K}_{\mathbf{af}}&\mathbf{K}_{\mathbf{aa}}\end{array}\right].$ (17)
Using properties of positive semi-definite matrices, we derive an upper bound $\widehat{\mathcal{U}}(M)$ to eq. (17):
$\mathcal{L}^{*}$ $\leq-\frac{(N_{n}+M_{\mathbf{a}})}{2}\log(2\pi)-\frac{1}{2}\log|\mathbf{Q}_{\mathbf{\hat{f}\hat{f}}}+\Sigma_{\mathbf{\hat{y}}}|-\frac{1}{2}\hat{\mathbf{y}}^{T}\left(\mathbf{Q}_{\mathbf{\hat{f}\hat{f}}}+t\mathrm{I}+\Sigma_{\mathbf{\hat{y}}}\right)^{-1}\hat{\mathbf{y}}+\Delta_{\mathbf{a}}\coloneqq\mathcal{\widehat{U}}(M),$
@@ 4.2 Approximation Quality Guarantees
Adding inducing points will eventually increase $\widehat{\mathcal{L}}$ until it reaches $\mathcal{L}^{*}$ [1, 25, 3]. If we add inducing points until $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})\leq\alpha$ we can guarantee the following:
@@
We cease the addition of points when ^$\widehat{\mathcal{L}}(M_{\mathbf{b}})>\widehat{\mathcal{U}}(M)-\alpha$. Given that $\widehat{\mathcal{U}}(M)\geq\mathcal{L}^{*}$, and assuming $\theta_{n}=\theta_{o}$, the rest follows from algebraic manipulation of eq. (4). The complete proof is provided in Appendix D. ∎
The first bound shows that if $\Psi$ is near zero, the KL to the true posterior is bounded by $\alpha$. While $\Psi$ depends on the true posterior and therefore cannot be computed, if the posterior in the previous iteration was exact, $\Psi$ would be equal to zero. The second bound gives us the insight that we are guaranteed to have our actual approximation $q_{n}(f)$ be within $\alpha$ nats of the best approximation that we can develop now, given the limitations of the approximations made in previous iterations.
@@ 4.3 Selecting a Threshold
In this final step of our online learning method, we must specify a heuristic for selecting $\alpha$ that does not require knowing any data in advance, while also working in a uniform way across datasets with different properties. A constant value for $\alpha$ does not work well, since the scale of the LML depends strongly on properties such as dataset size, and observation noise. This means that a tolerance of 1 nat [7] may be appropriate for a small dataset, but not for a large one.
We take inspiration from compression and MDL [18] and view the ELBO as being proportional to negative the code length that the model requires to encode the dataset. Instead of choosing a fixed $\alpha$, we select $\alpha$ such that we compress to within a small fraction of the optimal variational code. However, this is not well-defined, since it depends on the quantisation tolerance for our continuous variables. To avoid this, we take an independent random noise code as our baseline, and select $\alpha$ to be within some small fraction of the optimal variational code, relative to the random noise code, i.e. we want to guarantee:
$\alpha=\delta(\mathcal{L}^{*}-\mathcal{L}_{\text{noise}})\,,\qquad\mathcal{L}_{\text{noise}}=\sum_{n=1}^{N_{n}}\log\mathcal{N}(y_{n};\hat{\mu},\hat{\sigma}^{2})$ (20)
where $\hat{\mu}$ and $\hat{\sigma}^{2}$ are the average and variance of the observations for up to the current task and $\delta$ is a user-defined hyperparameter. The interpretation of $\delta$ allows it to be set in advance without needing prior data knowledge. For instance, to capture 95% of the information at each batch, $\delta$ should be set to 0.05. To remain computationally tractable for large batch sizes $N_{n}$, we can equivalently guarantee this by taking
$\alpha=\delta(\widehat{\mathcal{U}}(M)-\mathcal{L}_{\text{noise}})\;.$ (21)
There exist multiple options for selecting the number of inducing points to calculate $\widehat{\mathcal{U}}(M)$. When the batch size $N_{n}$ is small, including all new data points, i.e. setting $M=N_{n}+M_{\mathbf{a}}$, in the initial calculation of $\gamma$ is sufficient. Alternatively, the number of inducing points can be defined as the maximum between a predefined value and the size of $X_{n}$, for example, $M=\min(5000,N_{n}+M_{\mathbf{a}})$.
The algorithm for our inducing point selection method can be found in Appendix C. We name our approach Vegas Inducing Point Selection (VIPS), drawing an analogy to Las Vegas Algorithms [29]. These methods guarantee the accuracy of the output, however, their computational time fluctuates for every run [29].
@@ 5 Experiments
Throughout this section, we test the performance of our inducing point selection method VIPS across datasets with different properties. Our experiments aim to answer the following questions: How big should the size of a model be to effectively learn nearly all information within a dataset? How should the size of a model be dynamically adjusted in the context of continual learning? We assume a streaming scenario where the total number of points is unknown. In all experiments, the variational distribution and kernel hyperparameters are learnt by optimising the online lower bound in eq. (8). Additional experimental information is provided in Appendix F.

### v4

@@ 3.3 Online Sparse Gaussian Processes
^$\widehat{\mathcal{L}}=\mathbb{E}_{q_{n}(f)}\left[\log p(\mathbf{y}_{n}|f)\right]+\mathbb{E}_{q_{n}(f)}\left[\log l(\mathbf{a})\right]\\ -\operatorname{KL}\left[q_{n}(\mathbf{b})\mid\mid p(\mathbf{b}|\theta_{n})\right]-\log\mathcal{N}(\mathbf{a};0,\mathbf{K}^{\prime}_{\mathbf{aa}}+\mathbf{D}_{\mathbf{a}})\,,$ (7)
@@
Although $\widehat{\mathcal{L}}$ is a useful training objective for general likelihoods, the regression case we consider allows us to analytically find $q(\mathbf{b})$ (Bui et al., 2017) (see App. E), resulting in the lower bound
^$\widehat{\mathcal{L}}=\log\mathcal{N}\left(\hat{\mathbf{y}};\mathbf{0},\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b}\hat{\mathbf{f}}}+\Sigma_{\hat{\mathbf{y}}}\right)+\Delta\\ -\frac{1}{2}\operatorname{tr}\left[\mathbf{D}_{\mathbf{a}}^{-1}(\mathbf{K}_{\mathbf{aa}}-\mathbf{Q}_{\mathbf{aa}})\right]-\frac{1}{2\sigma^{2}}\operatorname{tr}(\mathbf{K}_{\mathbf{ff}}-\mathbf{Q}_{\mathbf{ff}})$ (8)
where $\mathbf{K}_{\hat{\mathbf{f}}\mathbf{b}}=\left[\mathbf{K}_{\mathbf{fb}}\quad\mathbf{K}_{\mathbf{ab}}\right]^{\top}$, $\Sigma_{\hat{\mathbf{y}}}=\text{diag}([\sigma_{y}^{2}\mathbf{I},\,\mathbf{D}_{\mathbf{a}}])$, $\hat{\mathbf{y}}=\left[\mathbf{y}_{n}\quad\mathbf{D}_{\mathbf{a}}\mathbf{S}{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}}\right]^{\top}$, and
$\Delta=-\frac{1}{2}\log\frac{|\mathbf{S}_{\mathbf{a}}|}{|\mathbf{K}_{\mathbf{aa}}^{\prime}||\mathbf{D}_{\mathbf{a}}|}+\frac{M_{\mathbf{a}}}{2}\log(2\pi)\\ -\frac{1}{2}\mathbf{m}_{\mathbf{a}}^{T}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}}+\frac{1}{2}\mathbf{m}_{\mathbf{a}}^{T}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{D}_{\mathbf{a}}\mathbf{S}_{\mathbf{a}}^{-1}\mathbf{m}_{\mathbf{a}},$
with $\mathbf{Q}_{\mathbf{ff}}=\mathbf{K}_{\mathbf{fb}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{bf}}$, $\mathbf{Q}_{\mathbf{aa}}=\mathbf{K}_{\mathbf{ab}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{ba}}$. All covariances are computed using the new hyperparameters $\theta_{n}$, except for $\mathbf{K}_{\mathbf{aa}}^{\prime}$ which is the covariance for the prior distribution $p(\mathbf{a}\mid\theta_{o})$. Finally, $M_{\mathbf{a}}=|\mathbf{a}|$ is the number of inducing points used at the previous batch. For calculating $\widehat{\mathcal{L}}$ in each batch, the computational complexity is $O(N_{n}M_{\mathbf{b}}^{2}+M_{\mathbf{b}}^{3})$ and the memory requirements are $O(M_{\mathbf{b}}^{2})$ where $M_{\mathbf{b}}$ is the total number of inducing points for the current batch.
To create a fully black-box solution, we still need to specify how to select the hyperparameters $\theta_{n}$, the number of inducing variables $M_{\mathbf{b}}$, and the inducing inputs $\mathbf{Z}_{n}$. We will always select $\theta_{n}$ by maximising $\widehat{\mathcal{L}}$ using L-BFGS. To select the locations $\mathbf{Z}_{n}$, we use the “greedy variance” criterion (Fine & Scheinberg, 2001; Foster et al., 2009; Burt et al., 2020). This leaves only the number of inducing variables $M_{\mathbf{b}}$ to be selected.
@@ 4 Automatically Adapting Approximation Capacity
We propose a method for adjusting the capacity of the approximation $M_{\bf b}$ to maintain accuracy. We propose to keep inducing points from old batches fixed, and select new inducing points from each incoming batch, with their locations set using the “greedy variance” criterion (Fine & Scheinberg, 2001; Foster et al., 2009; Burt et al., 2020). While optimising all inducing points does lead to a strictly better approximation, we avoid this for the sake of simplicity. The question remains: To achieve a level of accuracy, “how big is big enough?” To answer this, we consider the online ELBO as a function of the capacity ^$\widehat{\mathcal{L}}(M_{\bf b})$, and propose a threshold after which to stop adding new inducing variables.
@@ 4.1 Online Log Marginal Likelihood (LML) Upper Bound
The problem of selecting a sufficient number of inducing variables is also still open in the batch setting. One possible strategy is to use an upper bound to the marginal likelihood (Titsias, 2014) to bound $\mathrm{KL}[q(f)||p(f|\mathbf{y})]\leq\mathcal{U}-\mathcal{L}$, and stop adding inducing variables once this is below a tolerance $\alpha$. To extend this strategy to online learning, we begin by deriving an online upper bound, as a counterpart to the online ELBO from eq. (8). We follow the same strategy as Titsias (2014), by considering the highest possible value that our lower bound can attain. While in full-batch inference this is equal to the true LML, in our case this is obtained by keeping the inducing inputs from the previous iteration, and adding each new datapoint to the inducing set:
[…]
with $\quad\mathbf{K}_{\hat{\mathbf{f}}\hat{\mathbf{f}}}=\left[\begin{array}[]{cc}\mathbf{K}_{\mathbf{ff}}&\mathbf{K}_{\mathbf{fa}}\\ \mathbf{K}_{\mathbf{af}}&\mathbf{K}_{\mathbf{aa}}\end{array}\right].$ Using properties of positive semi-definite matrices, we derive an upper bound $\widehat{\mathcal{U}}(M)$ to eq. (9):
[…]
$\leq-\frac{(N+M_{\mathbf{a}})}{2}\log(2\pi)-\frac{1}{2}\log|\mathbf{K}_{\mathbf{\hat{f}b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b\hat{f}}}+\Sigma_{\mathbf{\hat{y}}}|$
$\quad-\frac{1}{2}\hat{\mathbf{y}}^{T}\left(\mathbf{K}_{\mathbf{\hat{f}b}}\mathbf{K}_{\mathbf{bb}}^{-1}\mathbf{K}_{\mathbf{b\hat{f}}}+t\mathbf{+}\Sigma_{\mathbf{\hat{y}}}\right)^{-1}\hat{\mathbf{y}}$
$\coloneqq\mathcal{\widehat{U}}(M),$
@@
Algorithm 1 Vegas Inducing Point Selection (VIPS)
  Output: Updated set $\mathbf{Z}_{n}=\mathbf{Z}_{o}\cup\{\mathbf{x}_{m^{\prime}}\}_{m^{\prime}=1}^{M^{\prime}}$, where $|\mathbf{Z}_{n}|=M_{\mathbf{b}}$.
  Initialise $\mathbf{Z}_{n}=\mathbf{Z}_{o}$.
  while $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})\leq\delta|\widehat{\mathcal{U}}(M)-\mathcal{L}_{noise}(\hat{\mu},\hat{\sigma})|$ do
  end while
@@ 4.2 Approximation Quality Guarantees
Adding inducing points will eventually increase $\widehat{\mathcal{L}}$ until it reaches $\mathcal{L}^{*}$ (Bauer et al., 2016; Matthews, 2017; Burt et al., 2020). If we add inducing points until $\widehat{\mathcal{U}}(M)-\widehat{\mathcal{L}}(M_{\mathbf{b}})\leq\alpha$ we can guarantee the following:
@@
We cease the addition of points when ^$\widehat{\mathcal{L}}(M_{\mathbf{b}})>\widehat{\mathcal{U}}(M)-\alpha$. Given that $\widehat{\mathcal{U}}(M)\geq\mathcal{L}^{*}$, and assuming $\theta_{n}=\theta_{o}$, the rest follows from algebraic manipulation of eq. (4). See App. C for the complete proof. ∎
The first bound shows that if $\Psi$ is near zero, the KL to the true posterior is bounded by $\alpha$. While $\Psi$ depends on the true posterior and therefore cannot be computed, if the posterior in the previous iteration was exact, $\Psi$ would be equal to zero. The second bound shows that we are guaranteed to have our actual approximation $q_{n}(f)$ be within $\alpha$ nats of the best approximation that we can develop, given the limitations of the approximations made in previous iterations.
@@ 4.3 Selecting a Threshold
In this final step of our online learning method, we must specify a heuristic for selecting $\alpha$ that does not require knowing any data in advance, while also working in a uniform way across datasets with different properties. A constant value for $\alpha$ does not work well, since the scale of the LML depends strongly on properties such as dataset size, and observation noise. This means that a tolerance of 1 nat (Cover, 1999) may be appropriate for a small dataset, but not for a large one.
As a principle for selecting the threshold, we take loose inspiration from compression and MDL (Grünwald & Roos, 2019), which takes the view of the ELBO being proportional to negative the code length that the model requires to encode the dataset. Intuitively, our desire to select an $\alpha$ such that our method captures a high proportion (e.g. 95%) of all the information in each batch, so that we can compress to within a small fraction of the optimal variational code. By itself, this is not well-defined since the code length depends on the quantisation tolerance for our continuous variables. To avoid this, we take an independent random noise code as our baseline, and select $\alpha$ to be within some small fraction of the optimal variational code, relative to the random noise code. We want to capture a high proportion of all the additional information that our model provides, relative to the noise model, i.e. we want our threshold to be:
$\alpha=\delta(\mathcal{L}^{*}-\mathcal{L}_{\text{noise}})\,,\quad\mathcal{L}_{\text{noise}}=\sum_{n=1}^{N_{n}}\log\mathcal{N}(y_{n};\hat{\mu},\hat{\sigma}^{2})$
where $\hat{\mu}$ and $\hat{\sigma}^{2}$ are the average and variance of the observations for up to the current task and $\delta$ is a user-defined hyperparameter. We validate that this approach leads to values of $\delta$ giving consistent behaviour across a wide range of datasets, which allows it to be set in advance without needing much prior knowledge of the dataset characteristics.
Calculating this threshold is intractable for large batch sizes $N_{n}$. However, if we change our stopping criterion to the more stringent upper bound
$\bar{\alpha}=\delta(\widehat{\mathcal{U}}(M)-\mathcal{L}_{\text{noise}})\,.$ (12)
and increase $M$ for calculating $\widehat{\mathcal{U}}$ as $M_{\mathbf{b}}$ is increased for calculating $\widehat{\mathcal{L}}$, we obtain the same guarantees as before but at a lower computational cost. However, this strategy is only worthwhile for very large batch sizes $N_{n}$, due to the importance of constant factors in the computational cost. In the common continual learning settings we investigate $N_{n}$ is small enough to allow computing $\mathcal{L}^{*}$.
The procedure for our inducing point selection method is detailed in Algorithm 1 and we give further details in App. B.1. We name our approach Vegas Inducing Point Selection (VIPS), drawing an analogy to Las Vegas Algorithms. These methods guarantee the accuracy of the output, however, their computational time fluctuates for every run (Motwani & Raghavan, 1995).
@@ 5 Experiments
We evaluate the performance of our adaptive inducing point selection method VIPS in a range of streaming scenarios where we assume the total number of observations is unknown. In all cases, the variational distribution, noise and kernel hyperparameters are optimised using the online lower bound (Eq. (8)). We provide further details in App. D.
Continual learning scenarios pose unique challenges. It is impossible to pre-determine memory allocation due to unknown input space coverage. Additionally, cross-validation for hyperparameter tuning is not feasible as it would require storing all data. Thus, an effective method must 1) have an adaptive memory that can grow with the demands of the data, 2) work with hyperparameters that can be set before training. The first part of our experiments demonstrates the benefits of using an adaptive model capacity in continual learning. The second part highlights the necessity of a single, pre-tunable hyperparameter that works across diverse settings. We benchmark VIPS against two other inducing point selection methods on streaming UCI datasets and a real-world robotics application.
5.1 Model size and data distribution
Figure 1 shows VIPS’s ability to adapt across three datasets with different characteristics, each divided into ten batches, illustrating how input distribution drives model growth as more data is seen. In the first dataset, each batch introduces data from new parts of the input space. Since each batch is equally novel, the model size grows linearly. In the second, the data remain within a fixed interval, leading to diminishing novelty in each batch, and a model size that converges to a fixed value. The third dataset mixes narrow-range uniform samples with occasional batches sampled from a heavy-tailed Cauchy distribution. This leads the model size to converge, with sporadic growth when novel data is observed.